OpenStudy (coconutjj):

A dart player throws a dart horizontally at a velocity of 12.4 m/s. The dart hits the board 0.32m below the height at which it is thrown. How far way is the player who threw the dart.

OpenStudy (anonymous):

will i get a medal if i anwser

OpenStudy (coconutjj):

yes if you can show your work clearly

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

12.72

OpenStudy (anonymous):

am i correct??

OpenStudy (coconutjj):

I need help, I don't know if your correct

OpenStudy (anonymous):

is it an optional question?

OpenStudy (coconutjj):

yes

OpenStudy (anonymous):

give me the options

OpenStudy (anonymous):

@CoconutJJ would you like me to help you?

OpenStudy (coconutjj):

YES!

OpenStudy (anonymous):

hey thats hurts

OpenStudy (anonymous):

@MinecrodMe you weren't offering a good way to solve it, that's why I butt in that's all...

OpenStudy (anonymous):

anyways so what we really need to do is solve for the amount of time it takes the dart to drop that much. Do you know how to do so?

OpenStudy (coconutjj):

0^2 = vi^2 + 2(-9.8)(0.32) I solved it like this for initial velocity

OpenStudy (coconutjj):

giving me vi = 2.5m/s

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

A hiker started her hike at 0.2 mi below sea level. At the end of her hike, she was 1.4 mi above sea level. What vertical distance did she hike? miles

OpenStudy (anonymous):

hmmm I think that you got the formula incorrect. What is the formula that you are using?

OpenStudy (coconutjj):

woops forgot 0.32^2

OpenStudy (coconutjj):

wait so its 1.416.. m/s

OpenStudy (anonymous):

hmmm I am still a bit puzzled at what you did so: You should have used this formula: \[x = x_{0}+v_{x}t+1/2(a)t^{2}\] did you do so?

OpenStudy (coconutjj):

yes...

OpenStudy (coconutjj):

solving for vi

OpenStudy (anonymous):

so what did you put in for each value?

OpenStudy (anonymous):

\[y = vt +\frac{ 1 }{ 2 }at ^{2}\] \[t = \sqrt{\frac{ 2y }{ g }}\] \[t = \sqrt{\frac{ 2(-0.32m) }{ -9.8m/s ^{2} }}= 0.25s\] \[x = vt + \frac{ 1 }{ 2 }at ^{2}\] \[x = vt +\frac{ 1 }{ 2 }(0)t ^{2}\] x = (12.4m/s)(0.25s) = 3.1m

OpenStudy (anonymous):

could you help me with my question??

OpenStudy (anonymous):

i will give a medal

OpenStudy (anonymous):

@MinecrodMe you need to open your own question, don't post your question on other people's post.

OpenStudy (anonymous):

@CoconutJJ so we don't need to solve for vx, since we actually know it

OpenStudy (anonymous):

what we need to solve for is t

OpenStudy (coconutjj):

vi = 12.4m/s

OpenStudy (anonymous):

oh I am sorry yes you are correct, we do know what vx is, (I actually wrote the wrong equation, this is the equation we need to use) \[y=y_{0}+v_{y}t+1/2(a)t^{2}\]

OpenStudy (anonymous):

Do you see what I did?

OpenStudy (anonymous):

or are you just extremely confused at this point?

OpenStudy (coconutjj):

yes

OpenStudy (anonymous):

ok I think that yes is meant for the first question I posed. Anyways can you tell me the values of these variables: 1) y 2) y0 3) vy 4) a ?

OpenStudy (coconutjj):

y = 0.32m

OpenStudy (coconutjj):

y0 = 0m

OpenStudy (coconutjj):

vy = 12.4m/s

OpenStudy (coconutjj):

a = 9.8m/s

OpenStudy (anonymous):

horizontally at a velocity of 12.4 m/s...vy = 0m/s

OpenStudy (anonymous):

ooh there's one thing that wrong. So vy doesn't equal 12.4m/s, vx is what equals 12.4m/s. We are told that it's the horizontal velocity.

OpenStudy (anonymous):

@starj9 gonna give u a medal for free :)

OpenStudy (coconutjj):

oh... I see..

OpenStudy (anonymous):

ok good! So now can you solve for t? aka how long it took for the dart to drop that far?

OpenStudy (coconutjj):

so 0t +1/2(9.8)(t) = 0.32?

OpenStudy (anonymous):

You know if yo looked up I already solved this step by step...

OpenStudy (anonymous):

you forgot the square on the t

OpenStudy (coconutjj):

oh yea..

OpenStudy (anonymous):

so what is t?

OpenStudy (coconutjj):

\[t=\sqrt{\frac{ 0.32 }{(0.5)(9.8) }}\]

OpenStudy (anonymous):

yes! ok so keep that value. Now think about this, the amount of time it took to drop that far, is the same amount of time it took the dart to go from your hand to the dart board. Does this make sense?

OpenStudy (coconutjj):

hmm... wait how?

OpenStudy (anonymous):

so when a dart leaves your hand it's both going forwards (due to it's initial velocity) and downwards (due to gravity). When it hits the dartboard, it stops, both dropping and going forward at the same time right?

OpenStudy (coconutjj):

yes..

OpenStudy (anonymous):

Because it's embedded into the dartboard. Now do you agree that the darts starts dropping and going forward at the same time?

OpenStudy (coconutjj):

yes

OpenStudy (anonymous):

do you agree that it stops going forward and downward at the same time?

OpenStudy (coconutjj):

yes

OpenStudy (anonymous):

so the time must be the same! So the amount of time it took to drop that far is the amount of time it took to travel that far to the dartboard. Make sense?

OpenStudy (coconutjj):

now it does..:)

OpenStudy (anonymous):

ok good. Now we know the time it took to travel that far. And we know the speed of the dart. So we use the same equation as above(but this time with x-components): \[x = x_{0}+v{x}t+1/2(a)t^{2}\] now tell me what variable we know and the values we put for them, and what we need to solve for.

OpenStudy (anonymous):

oops the equation should ne \[x= x_{0}+v_{x}t+1/2(a)t^{2}\]

OpenStudy (coconutjj):

vx = 12.4m/s t = 0.2555s a = 9.8 ??

OpenStudy (anonymous):

ahhh you are right to be a bit puzzled by the a, so the a as you probably know stands for acceleration. Now when we were looking at the dropping we knew that the dart was accelerating downwards due to gravity. But for moving horizontally does gravity affect you? aka does gravity pull/push you forwards or backwards?

OpenStudy (coconutjj):

ohh i see, so a=0?

OpenStudy (anonymous):

yes! There's really nothing that is pulling/pushing us forwards or backwards, and thus the acceleration in the horizontal plane is 0.

OpenStudy (jhannybean):

This question should be posted in the physics section..

OpenStudy (coconutjj):

so x = 12.4(0.2555) + 1/2(0)(0.2555^2) ?

OpenStudy (anonymous):

@CoconutJJ Make a new question for each of the questions you need help on!

OpenStudy (anonymous):

yup that's right!

OpenStudy (anonymous):

so how far are we from the dartboard?

OpenStudy (coconutjj):

3.1682m

OpenStudy (anonymous):

yup that looks about right!

OpenStudy (coconutjj):

THANK YOU @1DEA

OpenStudy (anonymous):

no problemo

OpenStudy (anonymous):

wow haha it censored me

OpenStudy (coconutjj):

THANK YOU TOO @StarJ9